Mike is x years old.
Sarah's age is Mike's age times 3 and plus 2.
Their combined age equals 50 years.
Set up an equation.

Answer by macston(5194)

S=Sarah's age; x=Mike's age
S+x=50 years
S=50yrs-x
S=3x+2
3x+2=50yrs-x
4x=48yrs
x=12yrs ANSWER 1: Mike is 12 years old.
S=50yrs-x=50yrs-12yrs=38yrs ANSWER 2:Sarah is 38 years old.
CHECK:
S=3x+2
38=3(12)+2
38=38